36.) Show a complete solution to the problem.
Two positive numbers differ by 11, and their square roots differ by 1. Find the numers.

Answer by scott8148(6628)

let x^2=larger number and y^2=smaller number __ x^2=y^2+11

x=y+1

substituting __ (y+1)^2=y^2+11 __ y^2+2y+1=y^2+11 __ 2y=10 __ y=5 __ y^2=25

substituting __ x^2=(25)+11 __ x^2=36